OpenStudy (anonymous):

A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation. Use any method to solve the equivalent system of equations (the new first equation with the original second equation). Prove that the solution for the equivalent system is the same as the solution for the original system of equations

OpenStudy (anonymous):

@Luigi0210 @wolfe8

OpenStudy (wolfe8):

Do the first part first. It's very direct.

OpenStudy (anonymous):

ok

OpenStudy (wolfe8):

And then this should help: http://www.purplemath.com/modules/systlin5.htm

OpenStudy (anonymous):

2 x 3 + 7 x 3 = 3 6x + 21y = 3 +-3x -4y +5 3x + 17y = 8

OpenStudy (wolfe8):

I see you are attempting elimination. What you want to do is multiply any of the equations with a number so that one of the unknowns from both equations are equal. That was you eliminate it.

OpenStudy (anonymous):

so i eliminate it ?

OpenStudy (anonymous):

if we eliminate, multiply the bottom by -2 6x + 21y = 3 -6x -34y = -16 ---------------- -13y = -13 y = 1 replace y by 1 6x + 21(1) = 3 solve for x 6x = -18 ; x = -3 so, is x=-3, and y=1 a solution to the original setup of 2x + 7y = 1 -3x – 4y = 5

OpenStudy (wolfe8):

If you did every step correctly, that should be right.